OpenStudy (anonymous):

Fan and medal :) Which equation represents the line that passes through the points (6,-3) and (-3,-9)?

OpenStudy (anonymous):

Please explain and help me through it

pooja195 (pooja195):

First you need to find the slope using the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

I'm thinking it's y+4=3/5(x+6)

OpenStudy (anonymous):

Wait so Y^2=6 y^1=-3/x^2=-4 - x^1=9?

OpenStudy (anonymous):

idk if that's right lol

pooja195 (pooja195):

Ok that's incorrect from what i can tell.... \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

What are y?

OpenStudy (anonymous):

Is that the right number (6,-3) (3)

pooja195 (pooja195):

Lable your x1 y1 and slope

OpenStudy (anonymous):

Mkay

OpenStudy (anonymous):

How do I do that :D

OpenStudy (anonymous):

I forget haven't done this in forever

pooja195 (pooja195):

What is the slope?

OpenStudy (anonymous):

I literally just got done with a 50 question star test my brain is fried

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{-9+3=?}{-3-6=?}=slope\]

OpenStudy (anonymous):

WHAT! ok, where's is the +3?

OpenStudy (anonymous):

Did you get that from another number that's isn't in the equation

pooja195 (pooja195):

when you have 2 negatives together they become positive \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{-9--3}{-3-6}=slope\] \[\LARGE \frac{-9+3=?}{-3-6=?}=slope\]

OpenStudy (anonymous):

-9+3=-6 -3-6 = -9

OpenStudy (anonymous):

Mkay I understand the negative together makes positive starting to understand now lol

pooja195 (pooja195):

Good now simplify that

OpenStudy (anonymous):

Wait what? how do I simplify that My gosh i'm bad at this

OpenStudy (anonymous):

I bet you're thinking i'm dumb right now huh?

OpenStudy (anonymous):

I'm literally like dead by laughing from this almost fell out of my chair haha

pooja195 (pooja195):



OpenStudy (anonymous):

Oh divide?

pooja195 (pooja195):

Pretty much now we have our slope m=-4/3 x1=6 y1=-3 Now... \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Plug in the numbers to get your answer

OpenStudy (anonymous):

-4/3 is 1.333

pooja195 (pooja195):

Done divide leave it as a fraction

pooja195 (pooja195):

I need to get to my next class @jabez177 or @KendrickLamar2014 can help u :)

pooja195 (pooja195):

Sorry!

OpenStudy (anonymous):

6-x^1=

OpenStudy (anonymous):

Wait class?

OpenStudy (anonymous):

Oh you're a teacher my bad ok sorry see yah

OpenStudy (anonymous):

Lamar

OpenStudy (anonymous):

Yo, What's happening?

OpenStudy (kendricklamar2014):

I have to go to. @jabez177 can probably help you

OpenStudy (anonymous):

Really dude

OpenStudy (anonymous):

Are you a teacher aswell?

OpenStudy (anonymous):

I got the answer y+3=3/5(x-6)

OpenStudy (anonymous):

I need help with one more please it's a graph and I'm bad with them so if anybody is there please help!

jabez177 (jabez177):

I have to go also... Sorry!!!

OpenStudy (anonymous):

Guys, Stop leaving and just help me with one question I know you have a little bit of time you can't be having to go it's literally 2:33 unless you started a class at 2:30 com'n Sorry just had to say it

jabez177 (jabez177):

I am so sorry!!! I had to go get a blood test so I didn't have any time at all. :P